OpenStudy (anonymous):

Write the equation of the quadratic function with roots -1 and 1 and a vertex at (0,-3)

OpenStudy (anonymous):

tmi

OpenStudy (anonymous):

write in vertex form?

OpenStudy (anonymous):

you have a choice of ways to do this, because you were given too much information method one you know it factors as \(a(x-1)(x+1)\) because you know the zeros you just don't know \(a\) but if you replace \(x\) by \(0\) and set the result equal to \(-3\) you will get it \[a(0-1)(0+1)=-3\] \[-a=-3\]\[a=3\]

OpenStudy (anonymous):

method 2 you are given the vertex as \((0.-3)\) so you know it has to look like \(y=ax^2-3\)

OpenStudy (anonymous):

this time find \(a\) since you are told if \(x=1\) then \(y=0\) solve \[a-3=0\] and so \(a=3\) and it is \(3x^2-3\) again

OpenStudy (anonymous):

okay now i need help with this one pleaseee Write the equation of the quadratic function with roots -10 and -8 and a vertex at (-9, -3).

OpenStudy (anonymous):

same idea exactly

OpenStudy (anonymous):

you want to use method one, or method two? makes no difference

OpenStudy (anonymous):

method 2

OpenStudy (anonymous):

vertex form is \(y=a(x-h)^2+k\) if the vertex is \((h,k)\) in your case the vertex is \((-9,-3)\) so your first step would be to write \[y=a(x+9)^2-3\]

OpenStudy (anonymous):

your second step would be to replace \(x\) by \(-8\) and set the result equal to zero and solve for \(a\) i.e. write \[a(-8+9)^2-3=0\] or \[a-3=0\] and now it should be easy

OpenStudy (anonymous):

okay thanks :)